# Problem List on Foliations and Diffeomorphism groups

### v1

@@ 2.1. Taut foliations and finite covers
Given a (transversely oriented) taut foliation $\mathscr{F}$ of a closed oriented 3-manifold, define the Euler class $e(\mathscr{F})\in H^{2}(M;\mathbb{R})$ as the Euler class of the oriented tangent plane field to $\mathscr{F}$. Thurston [Thu86] showed that taut foliations of closed hyperbolic 3-manifolds have Euler classes of dual Thurston norm at most one. Conversely, he conjectured that any even integral class $a\in H^{2}(M;\mathbb{R})$ is the Euler class of a taut foliation on $M$. This was negatively answered by Gabai and Yazdi [GY20, Yaz20]. The following is a virtual version of Thurston’s conjecture.
@@
Let $M$ be a closed hyperbolic 3-manifold whose first Betti number is at least one. Let $a\in H^{2}(M;\mathbb{R})$ be an (even) integral second cohomology class of dual Thurston norm one. Is there a finite cover $p\colon\tilde{M}\rightarrow M$ and a taut foliation $\mathscr{F}$ of $\tilde{M}$ such that the Euler class $e(\mathscr{F})$ of $\mathscr{F}$ is equal to $p^{*}(a)$? Here $p^{*}\colon H^{1}(M)\rightarrow H^{1}(\tilde{M})$, and we have used Poincaré duality to identify $H^{1}(M)\cong H^{2}(M)$ and $H^{1}(\tilde{M})\cong H^{2}(\tilde{M})$?
@@ 2.5. Engel structures
Now to something different (but related). Engel manifolds have canonical dynamics. Namely, there exists a unique line field $\mathscr{W}\subset\mathscr{D}$, called the kernel or characteristic line field, with the property $[\mathscr{W},\mathscr{E}]=\mathscr{E}$. We think of it as follows: $\mathscr{E}$ is invariant under the holonomy of $\mathscr{W}$ (so it can be seen as a contact structure in its leaf space), while $\mathscr{D}$ rotates along $\mathscr{W}$ (i.e. we imagine it as rotating within the leaf space contact structure).
@@ 2.6. Foliation by planes of $\mathbb{R}^{3}$
I would like to explain a little more. In the paper [Pal78] by Carlos Frederico Borges Palmeira, he claims that for foliations $\mathscr{F}$ and $\mathscr{F}^{\prime}$ by $(n-1)$-planes of simply-connected $n$-dimensional ($n\geqq 3$) manifolds $M$ and $M^{\prime}$, respectively, if the quotient non-Hausdorff 1-dimensional manifold $M/\mathscr{F}$ and $M^{\prime}/\mathscr{F}^{\prime}$ are diffeomorphic then there is a foliation preserving diffeomorphism $(M,\mathscr{F})$ and $(M^{\prime},\mathscr{F}^{\prime})$ inducing the diffeomorphism in the quotient. In the proof of this claim he uses his main lemma on page 118, which is the case of the theorem for saturated closed subsets $N\subset M$ and $N^{\prime}\subset M^{\prime}$ such that the interiors of $N/\mathscr{F}$ and $N^{\prime}/\mathscr{F}^{\prime}$ are diffeomorphic to $\mathbb{R}$. In the argument in the proof of the main lemma, he takes a family of disjoint neighborhoods of leaves of the boundary and asserts that this ensures the construction of foliation preserving diffeomorphism. In a part of my master thesis of 1978, On foliations of $\mathbb{R}^{n}$ whose leaves are homeomorphic to $\mathbb{R}^{n-1}$ [Tsu], I proved a weaker result only for $n=4$ or $n\geqq 6$. I had a reason to exclude $n=3$. That is because a locally supported infinite braid of the plane may not be extended to an isotopy of $\mathbb{R}^{2}$. Such an example is illustrated in Figures 5 and 6 of my thesis. It might be possible to argue without thinking about such isotopies. In higher dimensions, this braid question does not appear by a general-position argument.
@@ 3.1. Bounded cohomology of diffeomorphism groups
Let $\text{Diff}_{0}({\bf D}^{2},*)$ be the group of smooth orientation-preserving diffeomorphisms of ${\bf D}^{2}$ that fix a marked point * say at the origin. There are two bounded Euler classes on this group; one is induced by the restriction to the boundary and the other is by taking derivative at the origin. The difference between these two classes gives a nontrivial quasimorphism on $\text{Diff}_{0}({\bf D}^{2},*)$.
@@ 3.4. Milnor-Wood inequality, Euler class and Godbillon-Vey class
[…]
@@
The function $g_{\mathscr{E}(p)}(\epsilon)$, for $\epsilon>0$, is defined into the natural numbers union 00. For $\epsilon$ sufficiently small it obeys the upper bound:
@@
[…]
@@ 3.6. Amenable actions and planar groups and the Euler class
Under what condition on the holonomy of a compact leaf $L$ in a $C^{0}$ transversally oriented foliation of codimension $>1$, the leaf $L$ is stable? For example, suppose $\mathscr{F}$ is a $C^{0}$ codimension $2$ foliation and $L$ is a compact leaf whose normal microbundle has a trivial Euler class, its fundamental group is amenable and $H^{1}(L;\mathbb{R})=0$. Is $L$ stable?
@@ 3.7. Rigidity of group actions on the circle
Let $\Sigma$ be a closed Riemann surface of genus $>1$ and $\Gamma=\pi_{1}\Sigma$. Let $\rho$ be a $\Gamma$-action on ${\bf S}^{1}$. If $e(\rho)=e(\Sigma)$, then $\rho$ is semi-conjugate to the Fuchsian action $\rho_{0}$, i.e. $\exists\psi:{\bf S}^{1}\to{\bf S}^{1}$ s.t.  continuous of degree one, monotone and $(\rho,\rho_{0})$-equivariant. i.e.,
@@
Let $\Sigma$ be a closed Riemann surface of genus $>1$ and $\Gamma=\pi_{1}\Sigma$. Let $\rho$ be a geometric $\Gamma$-action on ${\bf S}^{1}$. Let $C_{\rho}$ be the connected component of $\operatorname{Hom}(\Gamma,\operatorname{Homeo}_{+}{\bf S}^{1})$ that contains $\rho$. Then every $\rho^{\prime}$ in $C_{\rho}$ is semi-conjugate to $\rho$.
@@ 4.1. Using commuting vector fields
Let S be a closed orientable surface of genus greater than $1$. Let $\rho:\pi_{1}(S)\to\mathrm{SL}(2,\mathbb{Z})$ be a representation. Let $\tilde{S}$ denote the universal cover of S and $\varphi:\pi_{1}(S)\to\text{Diff}(\tilde{S})$ be the natural action by deck transformations.
@@ 4.2. Using perturbations of group actions on the interval
Prove every leaf of a $C^{r}$-generic codimension $1$ oriented and co-oriented foliation on a $3$-manifold has genus $<2$.
@@
Consider four local $C^{r}$ diffeomorphisms $f_{a},f_{b},f_{c},f_{d}$ of $R$ whose definition domain contains $[-1,1]$ and having 00 as a common fixed point. Let $h=[f_{a},f_{b}][f_{c},f_{d}]$ be the product of the commutators.
Does there exist $\tilde{f}_{a},\tilde{f}_{b},\tilde{f}_{c},\tilde{f}_{d}$ coinciding with $f_{a},f_{b},f_{c},f_{d}$ out of $[-1,1]$, $C^{r}$-close to $f_{a},f_{b},f_{c},f_{d}$, so that
@@ 5.1. Topology of the space of commuting homeomorphisms
For commuting diffeomorphisms $f$ and $g$ of ${\bf S}^{1}$, to what extent we can say that the 2 cycle $(f,g)-(g,f)$ of $B\mathrm{Diff}({\bf S}^{1})^{\delta}$ is a boundary ?
@@ 7.2. Commutators of area preserving homeomorphisms of ${\bf S}^{2}$
What is $[G,G]$ What is the quotient $G/[G,G]$?
@@ 8.2. Exceptional minimal sets
The construction of exceptional minimal sets for codimension-$1$ foliations by Sacksteder [Sac64] and Rosenberg and Roussarie [RR70], along with Sacksteder’s theorem that the Denjoy minimal set does not exist for $C^{2}$-foliations [Sac65], can be viewed as the beginnings of the modern study of foliation dynamics. The works [CC88, CC02, hur, INA88, IM90, Mat88] all consider the structure of exceptional minimal sets in codimension-$1$. The more recent works of Deroin, Kleptsyn and Navas [DKN09, DFKN17] have made great strides towards understanding the dynamics of minimal sets in codimension 1, especially for the case of real analytic foliations.
@@
In codimension-$1$, there is a dichotomy for the derivative of the holonomy maps on a minimal set, in that they either have slow (subexponential) growth, or exponential growth [Sac64, hur, Hur11]. For higher codimensions, the transverse dynamics can be more complicated, as evidenced by the constructions of solenoidal minimal sets for flows of 3-manifolds, and hyperbolic minimal sets which have the transverse structure of a horseshoe. The work of Clark and Hurder [CH11] constructed solenoidal minimal sets for foliations with leaf dimensions at least 2, and this work has many references to prior works. A basic problem remains finding new constructions for such sets.
@@ 8.3. Powers of Euler class of flat circle bundles
$\chi_{\mathbb{Q}}^{k}\neq 0\in H^{2k}(B\textnormal{Diff}_{+}^{\omega}({\bf S}^{1})^{\delta};{\mathbb{Q}})$ for $k\geq 2$
@@ 8.4. A subgroup $G$ of $\mathrm{Diff}^{\omega}({\bf S}^{1})_{0}$
Can we show that $G$ is dense in $\mathrm{Diff}^{\omega}({\bf S}^{1})_{0}$
@@ 9. Homotopy theory of foliations
Thurston proved that $B\bar{\Gamma}_{q}$ is ($q+1$)-connected, and not ($12q+$ 1)-connected [Thu74a]. Haefliger and Bott proved that the continuous cohomology of $B\bar{\Gamma}_{q}$, in a certain sense, vanishes up to degree $2q$ [Hae79]. Question 9.1, which arose in 1972 at IAS Princeton during the collective discussions of the special year on Foliations [Law], was promoted to a conjecture by Thurston [Thu74a]; regarded as a question by Lawson [Law74]; while Haefliger wrote “one is tempted to conjecture that…” [Hae79].
@@
Consider the $\bar{\Gamma}_{q}$-structure $\gamma$ on the manifold $SO(q)$ whose microfoliation is the suspension of the canonical action of $SO(q)$ on $\mathbb{R}^{q}$. Is $\gamma$ null-homotopic in $B\bar{\Gamma}_{q}$
In other words, very concretely, is there a $\bar{\Gamma}_{q}$-structure $\bar{\gamma}$ on $SO(q)\times[0,1]$ whose restriction over $SO(q)\times 0$ is trivial (the microfoliation being the slice foliation of $SO(q)\times\mathbb{R}^{q}$ parallel to $SO(q)$) while the restriction of $\bar{\gamma}$ over $SO(q)\times 1$ is $\gamma$
@@
Here $B\Gamma^{+}_{q}$ is the classifying space for the Haefliger structures of codimension $q$ with arbitrary, oriented normal bundle. Hurder conjectured that (2) holds for every $q$; Vogt gave this conjecture the very concrete form of Question 9.2. The would-be corollary (3), by itself a classical question, would be a kind of reciproque to Bott’s obstructions, which state that the real cohomology algebra of $BSO(q)$ generated by the Pontrjagin classes vanishes in $B\Gamma^{+}_{q}$ in all degrees $>2q$.
@@
Question 9.3 (“Foliations with only one leaf”: Tsuboi [Tsu84b]).
Let $\Gamma_{q}^{+}$ be the groupoid of the germs of orientation-preserving local diffeomorphisms of $\mathbb{R}^{q}$, and $\Gamma_{q}^{0}\subset\Gamma_{q}^{+}$ be the subgroup fixing 00. Is $B\Gamma_{q}^{0}$ homotopically trivial in $B\Gamma_{q}^{+}$
This amounts to ask if any (orientable) Haefliger structure whose microfoliation has the zero section as a leaf is homotopically trivial. Tsuboi answered positively for $q=1$.
@@
i) Is every codimension-$q$ Riemannian foliation homotopically trivial in $B\Gamma_{q}$
ii) Is every codimension-$q$ Lie foliation homotopically trivial in $B\bar{\Gamma}_{q}$
@@ 9.3.1. Yoshihiko Mitsumatsu
For $r=\omega$ $q\geq 2$ not very much is known. For $q=1$, Haefliger proved that $B\overline{\Gamma}^{\omega}_{1}$ is a $K(\pi,1)$ space for a big discrete group $\pi=\Gamma_{H}$ which we call the Haefliger group. He also showed that $\Gamma_{H}$ is perfect, continuously generated, and of torsion free. Therefore we have $H_{1}(B\overline{\Gamma}^{\omega}_{1};{\mathbb{Z}})=H_{1}(\Gamma_{H};{\mathbb{Z}})=0$. As the Thurston (or Bott-Brooks) cycles show that the Godbillon-Vey invariant gives a surjection $H_{3}(B\Gamma_{H};{\mathbb{Z}})\to{\mathbb{R}}$ which factors through $H_{3}(B\overline{\Gamma}^{\infty}_{1};{\mathbb{Z}})=\pi_{1}(B\overline{\Gamma}^{\infty}_{1})$. Tsuboi [Tsu84a] constructed a homomorphic section ${\mathbb{R}}\to\pi_{1}(B\overline{\Gamma}^{\infty}_{1})$ and recently Morita (et al.) modified it into ${\mathbb{R}}\to H_{3}(B\Gamma_{H};{\mathbb{Z}})$. So we have $\pi_{1}(B\overline{\Gamma}^{\infty}_{1})\cong{\mathbb{R}}\oplus\mathrm{ker}GV$ as well as $H_{3}(B\Gamma_{H};{\mathbb{Z}})\cong{\mathbb{R}}\oplus\mathrm{ker}GV$.
@@
As $B\Gamma_{H}$ has no $\pi_{2}$ the 2-sphere does not support $H_{2}(B\Gamma_{H};{\mathbb{Z}})$. It is also known that 2-torus $T^{2}$ does not support $H_{2}(B\Gamma_{H};{\mathbb{Z}})$.
@@
(Tsuboi 2017) Determine $H_{2}(B\Gamma_{H};{\mathbb{Z}})=0$.
If a loop in $B\Gamma_{H}$ is free homotopic to one which is transverse to the $\Gamma^{\omega}_{1}$ structure, it presents a non trivial element in $\pi_{1}(B\overline{\Gamma}^{\omega}_{1})=\Gamma_{H}$ which is of order infinite.

### v2

@@ 2.1. Taut foliations and finite covers
Given a (transversely oriented) taut foliation $\mathscr{F}$ of a closed oriented 3-manifold, define the Euler class $e(\mathscr{F})\in H^{2}(M;\mathbb{R})$ as the Euler class of the oriented tangent plane field to $\mathscr{F}$. Thurston [Thu86] showed that taut foliations of closed hyperbolic 3-manifolds have Euler classes of dual Thurston norm at most one. Conversely, he conjectured that any even integral class $a\in H^{2}(M;\mathbb{R})$ of dual norm one is the Euler class of a taut foliation on $M$. This was negatively answered by Gabai and Yazdi [GY20, Yaz20]. The following is a virtual version of Thurston’s conjecture.
@@
[…]
@@ 2.5. Engel structures
Now to something different (but related). Engel manifolds have canonical dynamics. Namely, there exists a unique line field $\mathscr{W}\subset\mathscr{D}$, called the kernel or characteristic line field, with the property $\mathscr{W},\mathscr{E}]=\mathscr{E}$. We think of it as follows: $\mathscr{E}$ is invariant under the holonomy of $\mathscr{W}$ (so it can be seen as a contact structure in its leaf space), while $\mathscr{D}$ rotates along $\mathscr{W}$ (i.e. we imagine it as rotating within the leaf space contact structure).
@@ 2.6. Foliation by planes of $\mathbb{R}^{3}$
I would like to explain a little more. In the paper [Pal78] by Carlos Frederico Borges Palmeira, he claims that for foliations $\mathscr{F}$ and $\mathscr{F}^{\prime}$ by (n1)(n-1)-planes of simply-connected $n$-dimensional ($n\geqq 3$) manifolds $M$ and $M^{\prime}$, respectively, if the quotient non-Hausdorff 1-dimensional manifold $M/\mathscr{F}$ and $M^{\prime}/\mathscr{F}^{\prime}$ are diffeomorphic then there is a foliation preserving diffeomorphism $(M,\mathscr{F})$ and $(M^{\prime},\mathscr{F}^{\prime})$ inducing the diffeomorphism in the quotient. In the proof of this claim he uses his main lemma on page 118, which is the case of the theorem for saturated closed subsets $N\subset M$ and $N^{\prime}\subset M^{\prime}$ such that the interiors of $N/\mathscr{F}$ and $N^{\prime}/\mathscr{F}^{\prime}$ are diffeomorphic to $\mathbb{R}$. In the argument in the proof of the main lemma, he takes a family of disjoint neighborhoods of leaves of the boundary and asserts that this ensures the construction of foliation preserving diffeomorphism. In a part of my master thesis of 1978, On foliations of $\mathbb{R}^{n}$ whose leaves are homeomorphic to $\mathbb{R}^{n-1}$ [Tsu], I proved a weaker result only for $n=4$ or $n\geqq 6$. I had a reason to exclude $n=3$. That is because a locally supported infinite braid of the plane may not be extended to an isotopy of $\mathbb{R}^{2}$. Such an example is illustrated in Figures 5 and 6 of my thesis. It might be possible to argue without thinking about such isotopies. In higher dimensions, this braid question does not appear by a general-position argument.
@@ 3.1. Bounded cohomology of diffeomorphism groups
Let $\text{Diff}_{0}({\bf D}^{2},*)$ be the group of smooth orientation-preserving diffeomorphisms of ${\bf D}^{2}$ that fix a marked point * say at the origin. There are two bounded Euler classes on this group; one is induced by the restriction to the boundary and the other is by taking derivative at the origin. The difference between these two classes gives a non-trivial quasimorphism on $\text{Diff}_{0}({\bf D}^{2},*)$.
@@ 3.4. Milnor-Wood inequality, Euler class and Godbillon-Vey class
The boundedness of the Euler class for flat ${\bf S}^{1}$-bundles is known as the Milnor-Wood inequality [Mil58], [Woo71]. For a $C^{r}$-flat ${\bf S}^{1}$-bundle $E\xrightarrow{p}\Sigma_{g}$ over a closed oriented surface $\Sigma_{g}$ of genus $g>0$ states
@@
The function $g_{\mathscr{E}(p)}(\epsilon)$, for $\epsilon>0$, is defined into the natural numbers union 0. For $\epsilon$ sufficiently small it obeys the upper bound:
@@
Suppose we are given an ${\bf S}^{3}$-bundle ${\bf S}^{3}\to E\xrightarrow{p}M$ over a $4$-manifold $M$ with nontrivial Euler class and $p_{1}$. Mather-Thurston’s theory says we can find a bordism of this bundle to a $C^{0}$-flat bundle ${\bf S}^{3}\to E^{\prime}\to M^{\prime}$. Can one bound the simplicial volume of $M^{\prime}$ so that it does not depend on the bundle map $p$?
@@ 3.6. Amenable actions and planar groups and the Euler class
Under what condition on the holonomy of a compact leaf $L$ in a $C^{0}$ transversally oriented foliation of codimension >1>1, the leaf $L$ is stable? For example, suppose $\mathscr{F}$ is a $C^{0}$ codimension $2$ foliation and $L$ is a compact leaf whose normal microbundle has a trivial Euler class, its fundamental group is amenable and $H^{1}(L;\mathbb{R})=0$. Is $L$ stable?
@@ 3.7. Rigidity of group actions on the circle
Let $\Sigma$ be a closed Riemann surface of genus >1>1 and $\Gamma=\pi_{1}\Sigma$. Let $\rho$ be a $\Gamma$-action on ${\bf S}^{1}$. If $e(\rho)=e(\Sigma)$, then $\rho$ is semi-conjugate to the Fuchsian action $\rho_{0}$, i.e. $\exists\psi:{\bf S}^{1}\to{\bf S}^{1}$ s.t.  continuous of degree one, monotone and $(\rho,\rho_{0})$-equivariant. i.e.,
@@
Let $\Sigma$ be a closed Riemann surface of genus >1>1 and $\Gamma=\pi_{1}\Sigma$. Let $\rho$ be a geometric $\Gamma$-action on ${\bf S}^{1}$. Let $C_{\rho}$ be the connected component of $\operatorname{Hom}(\Gamma,\operatorname{Homeo}_{+}{\bf S}^{1})$ that contains $\rho$. Then every $\rho^{\prime}$ in $C_{\rho}$ is semi-conjugate to $\rho$.
@@ 4.1. Using commuting vector fields
[…]
@@ 4.2. Using perturbations of group actions on the interval
Prove every leaf of a $C^{r}$-generic codimension $1$ oriented and co-oriented foliation on a $3$-manifold has genus <2<2.
@@
Consider four local $C^{r}$ diffeomorphisms $f_{a},f_{b},f_{c},f_{d}$ of $R$ whose definition domain contains [1,1][-1,1] and having 0 as a common fixed point. Let $h=[f_{a},f_{b}][f_{c},f_{d}]$ be the product of the commutators.
Does there exist $\tilde{f}_{a},\tilde{f}_{b},\tilde{f}_{c},\tilde{f}_{d}$ coinciding with $f_{a},f_{b},f_{c},f_{d}$ out of [1,1][-1,1], $C^{r}$-close to $f_{a},f_{b},f_{c},f_{d}$, so that
@@ 5.1. Topology of the space of commuting homeomorphisms
For commuting diffeomorphisms $f$ and $g$ of ${\bf S}^{1}$, to what extent we can say that the 2 cycle (f,g)(g,f)(f,g)-(g,f) of $B\mathrm{Diff}({\bf S}^{1})^{\delta}$ is a boundary ?
@@ 7.2. Commutators of area preserving homeomorphisms of ${\bf S}^{2}$
What is $[G,G]$ ? What is the quotient $G/[G,G]$?
@@ 8.2. Exceptional minimal sets
The construction of exceptional minimal sets for codimension-$1$ foliations by Sacksteder [Sac64] and Rosenberg and Roussarie [RR70], along with Sacksteder’s theorem that the Denjoy minimal set does not exist for $C^{2}$-foliations [Sac65], can be viewed as the beginnings of the modern study of foliation dynamics. The works [CC88, CC02, Hur91, INA88, IM90, Mat88] all consider the structure of exceptional minimal sets in codimension-$1$. The more recent works of Deroin, Kleptsyn and Navas [DKN09, DFKN17] have made great strides towards understanding the dynamics of minimal sets in codimension 1, especially for the case of real analytic foliations.
@@
In codimension-$1$, there is a dichotomy for the derivative of the holonomy maps on a minimal set, in that they either have slow (subexponential) growth, or exponential growth [Sac64, Hur91, Hur11]. For higher codimensions, the transverse dynamics can be more complicated, as evidenced by the constructions of solenoidal minimal sets for flows of 3-manifolds, and hyperbolic minimal sets which have the transverse structure of a horseshoe. The work of Clark and Hurder [CH11] constructed solenoidal minimal sets for foliations with leaf dimensions at least 2, and this work has many references to prior works. A basic problem remains finding new constructions for such sets.
@@ 8.3. Powers of Euler class of flat circle bundles
$\chi_{\mathbb{Q}}^{k}\neq 0\in H^{2k}(B\textnormal{Diff}_{+}^{\omega}({\bf S}^{1})^{\delta};{\mathbb{Q}})$ for $k\geq 2$ ?
@@ 8.4. A subgroup $G$ of $\mathrm{Diff}^{\omega}({\bf S}^{1})_{0}$
Can we show that $G$ is dense in $\mathrm{Diff}^{\omega}({\bf S}^{1})_{0}$ ?
@@ 9. Homotopy theory of foliations
Thurston proved that $B\bar{\Gamma}_{q}$ is ($q+1$)-connected, and not ($2q+1$)-connected [Thu74a]. Haefliger and Bott proved that the continuous cohomology of $B\bar{\Gamma}_{q}$, in a certain sense, vanishes up to degree $2q$ [Hae79]. Question 9.1, which arose in 1972 at IAS Princeton during the collective discussions of the special year on Foliations [Law], was promoted to a conjecture by Thurston [Thu74a]; regarded as a question by Lawson [Law74]; while Haefliger wrote “one is tempted to conjecture that…” [Hae79].
@@
Consider the $\bar{\Gamma}_{q}$-structure $\gamma$ on the manifold $SO(q)$ whose microfoliation is the suspension of the canonical action of $SO(q)$ on $\mathbb{R}^{q}$. Is $\gamma$ null-homotopic in $B\bar{\Gamma}_{q}$ ?
In other words, very concretely, is there a $\bar{\Gamma}_{q}$-structure $\bar{\gamma}$ on $SO(q)\times[0,1]$ whose restriction over $SO(q)\times 0$ is trivial (the microfoliation being the slice foliation of $SO(q)\times\mathbb{R}^{q}$ parallel to $SO(q)$) while the restriction of $\bar{\gamma}$ over $SO(q)\times 1$ is $\gamma$ ?
@@
Here $B\Gamma^{+}_{q}$ is the classifying space for the Haefliger structures of codimension $q$ with arbitrary, oriented normal bundle. Hurder conjectured that (2) holds for every $q$; Vogt gave this conjecture the very concrete form of Question 9.2. The would-be corollary (3), by itself a classical question, would be a kind of reciproque to Bott’s obstructions, which state that the real cohomology algebra of $BSO(q)$ generated by the Pontrjagin classes vanishes in $B\Gamma^{+}_{q}$ in all degrees >2q>2q.
@@
Question 9.3 (“Foliations with only one leaf”: see Tsuboi [Tsu84b]).
Let $\Gamma_{q}$ be the groupoid of the germs of local diffeomorphisms of $\mathbb{R}^{q}$, and $G_{q}\subset\Gamma_{q}$ be the subgroup fixing 0. Let $B\bar{G}_{q}$ be the homotopy-theoretic fibre of the map $BG_{q}\to BGL_{q}(\mathbb{R})$ classifying the normal bundle of the universal $G_{q}$-structure. Is $B\bar{G}_{q}$ homotopically trivial in $B\bar{\Gamma}_{q}$ ?
This amounts to asking if any (framed) Haefliger structure whose microfoliation has the zero section as a leaf is homotopically trivial. Tsuboi answered positively for $q=1$.
@@
i) Is every codimension-$q$ Riemannian foliation homotopically trivial in $B\Gamma_{q}$ ?
ii) Is every codimension-$q$ Lie foliation homotopically trivial in $B\bar{\Gamma}_{q}$ ?
@@ 9.3.1. Yoshihiko Mitsumatsu
For $r=\omega$ $q\geq 2$ not very much is known. For $q=1$, Haefliger proved that $B\overline{\Gamma}^{\omega}_{1}$ is a $K(\pi,1)$ space for a big discrete group $\pi=\Gamma_{H}$ which we call the Haefliger group. He also showed that $\Gamma_{H}$ is perfect, continuously generated, and torsion-free. Therefore we have $H_{1}(B\overline{\Gamma}^{\omega}_{1};{\mathbb{Z}})=H_{1}(\Gamma_{H};{\mathbb{Z}})=0$. As the Thurston (or Bott-Brooks) cycles show that the Godbillon-Vey invariant gives a surjection $H_{3}(B\Gamma_{H};{\mathbb{Z}})\to{\mathbb{R}}$ which factors through $H_{3}(B\overline{\Gamma}^{\infty}_{1};{\mathbb{Z}})=\pi_{3}(B\overline{\Gamma}^{\infty}_{1})$. Tsuboi [Tsu84a] constructed a homomorphic section ${\mathbb{R}}\to\pi_{3}(B\overline{\Gamma}^{\infty}_{1})$ and recently Morita (et al.) modified it into ${\mathbb{R}}\to H_{3}(B\Gamma_{H};{\mathbb{Z}})$. So we have $\pi_{3}(B\overline{\Gamma}^{\infty}_{1})\cong{\mathbb{R}}\oplus\mathrm{ker}GV$ as well as $H_{3}(B\Gamma_{H};{\mathbb{Z}})\cong{\mathbb{R}}\oplus\mathrm{ker}GV$.
@@
As $B\Gamma_{H}$ has a trivial $\pi_{2}$ there is no nontrivial spherical cycle in $H_{2}(B\Gamma_{H};{\mathbb{Z}})$. It is also known that all maps from the 2-torus $T^{2}$ to $B\Gamma_{H}$ yield trivial cycles in $H_{2}(B\Gamma_{H};{\mathbb{Z}})$.
@@
(Tsuboi [Tsu17]) Determine whether $H_{2}(B\Gamma_{H};{\mathbb{Z}})=0$.
If a loop in $B\Gamma_{H}$ is freely homotopic to one that is transverse to the $\Gamma^{\omega}_{1}$ structure, it presents a non trivial element in $\pi_{1}(B\overline{\Gamma}^{\omega}_{1})=\Gamma_{H}$ which is of infinite order.